# Real-IAD Variety: Pushing Industrial Anomaly Detection Dataset to a Modern Era

### v1

@@ 3.3 Dataset Characteristics and Statistical Analysis
Datasets Classes The Number of Images
Image
Resolution
Anomaly
Masks
Multiple
Views
Defect Types
@@
Statistical Distribution Analysis. Figure 3 illustrates the advantages and challenges of Real-IAD Variety compared to existing IAD datasets across multiple dimensions: (a) Anomalous region proportion: Compared to Real-IAD [17], Real-IAD Variety exhibits a broader and more uniform distribution of anomaly region proportions (Figure 3a), encompassing defects of varying sizes and thereby increasing detection difficulty. (b) Defect aspect ratio: The defect aspect ratio distribution in Real-IAD Variety is comparably diverse to Real-IAD (Figure 3b), presenting substantial challenges for models to discriminate defects of various geometric configurations. (c) Material diversity: Real-IAD Variety encompasses 24 material types (Figure 3c) across 28 industrial domains (Figure 1), providing comprehensive coverage from a practical application perspective. (d) Color spectrum: The dataset captures a wide color distribution (Figure 3d), which is critical for evaluating color-based anomaly detection approaches. These multifaceted characteristics substantially enhance dataset diversity and provide robust support for large-scale anomaly detection algorithm research.
Representative Defect Visualization. Figure 2d presents qualitative visualizations of 23 distinct defect types across various object categories. Notably, identical defect types manifest in diverse visual forms across different materials, reflecting the complexity of real-world industrial scenarios. These defect types, derived from authentic production line observations, enhance the practical applicability of Real-IAD Variety.
@@ 4.1 Dataset Protocol and Evaluation Metrics
Training and Testing Protocol. The Real-IAD Variety dataset is partitioned into training and testing subsets, comprising 3,991 normal samples (19,955 images) for training and 35,799 samples (3,990 normal and 31,809 anomalous, totaling 178,995 images) for testing. Notably, the testing subset exhibits a nearly balanced distribution between normal and anomalous instances. To systematically evaluate the impact of category count on model performance, we partition Real-IAD Variety into three subsets: S1, S2, and S3, containing 30, 60, and 100 categories, respectively. The category selection employs randomization to mitigate potential biases in color and material representation.
Evaluation Metrics. Following established anomaly detection protocols, we employ three standard metrics: Image-level Area Under the Receiver Operating Characteristic Curve (I-AUROC) [27], Pixel-level AUROC (P-AUROC) [27], and Pixel-level Area Under the Precision-Recall Curve (P-AUPR) [15] to assess method effectiveness across classification and localization tasks.
@@ 4.2 Benchmark on Multi-Class Unsupervised AD
Table 2: MUAD performance comparisons on Real-IAD Variety dataset. Dis., Emb., and Rec., respectively, represent discrimination-based, embedding-based, and reconstruction-based methods. Bold indicates the best results.
Methods I-AUROC P-AUROC P-AUPR
Dis.
DREAM [27] 49.6 52.1 1.3
SimpleNet [11] 54.5 70.6 3.6
Emb.
CFA [10] 52.3 56.8 2.2
CFLOW-AD [33] 69.8 86.5 18.0
Rec.
RD [30] 73.0 89.5 18.9
MabaAD [12] 79.5 90.8 28.2
UniAD [4] 64.2 86.4 11.3
DesTSeg [34] 76.2 67.4 35.2
Dinomaly [6] 81.4 91.5 37.6
Dinomaly+ 81.6 91.9 40.7
Results and Analysis under MUAD.
Table 3: MUAD performance comparisons across different anomaly detection datasets. Performance is measured using I-AUROC, P-AUROC, and P-AUPR metrics.
Datasets # Classes CFLOW-AD [33] SimpleNet [11] UniAD [4] MambaAD [12] Dinomaly [6] Dinomaly+
MVTec 15 80.4 90.7 37.1 78.2 81.0 24.8 96.5 96.8 44.7 98.6 97.7 56.3 99.6 98.4 69.3 99.7 98.7 76.9
ViSA 12 69.0 91.4 16.8 89.2 95.3 33.1 90.8 98.4 33.6 94.3 98.5 39.4 98.7 98.7 53.2 98.8 98.7 54.6
Real-IAD 30 55.7 81.3 1.6 57.2 75.7 2.8 83.0 97.3 21.1 86.3 98.5 33.0 89.3 98.8 42.8 90.0 98.9 47.6
Real-IAD Variety S1 30 55.0 63.8 2.9 75.0 84.8 21.7 79.9 89.7 24.4 87.0 92.3 39.4 90.1 92.4 54.1 90.5 92.5 57.0
Real-IAD Variety S2 60 54.2 60.0 2.6 67.1 82.0 13.3 66.9 86.7 14.1 85.8 92.3 37.8 88.1 92.4 49.0 88.7 92.7 53.1
Real-IAD Variety S3 100 52.9 57.6 2.4 59.3 74.7 5.7 66.7 86.9 12.9 85.0 92.3 35.7 85.4 92.2 43.8 85.7 92.3 47.2
Real-IAD Variety 160 52.3 56.8 2.2 54.5 70.6 3.6 64.2 86.4 11.3 79.5 90.8 28.2 81.4 91.5 37.6 81.6 91.7 42.4
@@
Unless otherwise specified, MUAD models are trained at $\times256$ and a total of 100 epochs. For Dinomaly [6], we follow the original setting with $\times448$ resolution for 50,000 steps (e.g., approximately 160 epochs for Real-IAD Variety). Dinomaly+ extends Dinomaly [6] by incorporating an additional refined segmentation head [5], trained for 4 epochs on pseudo-anomalous data to enable coarse-to-fine anomaly localization.
@@
Competitive Methods. We evaluate state-of-the-art unsupervised anomaly detection methods to comprehensively assess their scalability to large-scale datasets. Specifically, we benchmark discrimination-based methods (DRAEM [27], SimpleNet [11]), embedding-based methods (CFA [10], CFLOW-AD [33]), and reconstruction-based methods (RD [30], MambaAD [12], UniAD [4], DesTSeg [34], Dinomaly [6], and Dinomaly+) on Real-IAD Variety. Considering computational resource constraints, we focus on representative and strong baselines—CFLOW-AD [33], SimpleNet [11], UniAD [4], MambaAD [12], Dinomaly [6], and Dinomaly+ for comprehensive MUAD evaluation. Notably, PatchCore [31] is excluded from this comparison due to its memory bank mechanism, which incurs prohibitive memory consumption as category count increases, leading to potential out-of-memory issues on large-scale datasets like Real-IAD Variety.
Table 2 presents quantitative results of different methods on Real-IAD Variety. Several critical observations emerge that differ from findings on small-scale datasets:
All evaluated methods experience substantial and non-saturating performance degradation on Real-IAD Variety, indicating the dataset’s significant challenge for future research and confirming successful avoidance of the metric saturation phenomenon frequently observed in smaller-scale IAD benchmarks.
Discrimination-based approaches (DRAEM [27], SimpleNet [11]) and the embedding-based method CFA [10] exhibit near-failure performance, with I-AUROC scores approaching the 50% random guessing threshold. This demonstrates their severe lack of generalization capability when confronted with large-scale, high-variance normal data distributions.
Reconstruction-based methods maintain relatively satisfactory performance despite the challenging conditions. Notably, Dinomaly+ (building upon Dinomaly [6] and OneNIP [5]) achieves the best overall results across all metrics (I-AUROC: 81.6%, P-AUROC: 91.9%, P-AUPR: 40.7%), underscoring its superior generalization and robustness across large-scale multi-class scenarios. In contrast, UniAD [4], an early MUAD method, exhibits one of the most severe performance drops among reconstruction-based approaches.
@@
Performance Trends with Increasing Categories. To evaluate performance trends as category count increases, we selected five representative methods across different paradigms. Table 3 presents results on several datasets and Real-IAD Variety with incrementally increasing categories: S1 (30), S2 (60), S3 (100), and the full 160 categories (visualized in Figure 4). Several consistent conclusions emerge under the MUAD setting:
Real-IAD Variety presents greater challenges compared to the original Real-IAD dataset. Despite maintaining equivalent category counts (30 classes), performance metrics on Real-IAD Variety are consistently lower, suggesting higher inherent complexity or greater intra-class variance.
A fundamental scalability challenge of MUAD is revealed: consistent performance degradation occurs across all evaluated methods as category count increases. Crucially, the magnitude of degradation is method-dependent, highlighting that different architectures possess varying levels of category generalization capability and robustness to scale-up.
@@
Methods leveraging powerful backbones (e.g., DINOv2-R [41]) and advanced architectures (e.g., Transformer decoders), such as Dinomaly and Dinomaly+, consistently achieve substantial performance superiority across all datasets. This advantage is especially prominent on high-scale benchmarks, underscoring their effectiveness over normalizing flow-based (e.g., CFLOW-AD) or simpler backbone-based models (e.g., SimpleNet, UniAD, MambaAD).
Table 4: Comparisons of ZSAD and FSAD methods across different benchmarks. Performance is measured using I-AUROC, P-AUROC, and P-AUPR metrics. Delta indicates the performance gap between maximum and minimum values.
Datasets # Classes Zero-Shot 1-Shot 2-Shot
AnomalyCLIP [7] AdaCLIP [38] VCPCLIP [39] AdaptCLIP [8] AdaptCLIP [8] AdaptCLIP [8]
Real-IAD 30 69.5 94.8 26.6 71.8 80.3 34.6 71.7 95.9 29.2 74.2 94.9 28.2 81.8 97.1 36.6 82.9 97.3 37.8
Real-IAD Variety S1 30 67.7 87.9 33.7 73.5 73.4 34.8 74.0 88.6 33.6 73.2 88.1 36.7 79.4 90.7 49.0 81.2 91.0 50.4
Real-IAD Variety S2 60 69.0 88.2 33.5 74.0 74.3 37.7 76.1 89.1 36.0 72.6 88.1 36.5 78.7 91.0 48.0 80.4 91.3 49.8
Real-IAD Variety S3 100 70.0 89.1 33.8 75.2 75.1 38.5 76.1 90.2 36.5 72.8 89.1 36.8 79.2 91.8 48.3 81.0 92.0 50.2
Real-IAD Variety 160 69.8 89.3 32.9 74.4 74.5 38.9 75.9 90.6 36.2 73.0 89.2 36.2 79.7 92.0 48.0 81.4 92.2 49.7
Delta   $\downarrow$ 2.3 6.9 7.2 3.4 7.0 4.3 4.4 7.3 7.3 1.6 6.8 8.6 3.1 6.4 12.4 2.5 6.3 12.6

### v2

@@ 3.2 Comparison with Existing IAD Datasets
Datasets Classes Number of Images Image Resolution Anomaly Masks Multiple Views Defect Types
@@
Kaputt [19] 48,376 70,951 29,316 100,267 $\sim2,048$ 7
[…]
Figure 5: The distribution of normal and abnormal images (first row), defect spatial locations (second row), defects scales (third row) and defect area ratios (last row) in our Real-IAD Variety dataset.
3.3 Dataset Characteristics and Statistical Analysis
Statistical Distribution Analysis. To provide a comprehensive understanding of Real-IAD Variety, we perform a multi-dimensional statistical analysis as illustrated in Fig. 4 and Fig. 5: (a) Material Diversity: Real-IAD Variety encompasses 24 material types (Fig. 4c) across 28 industrial domains (Fig. 2), providing comprehensive coverage from a practical application perspective. (b) Color Spectrum: The dataset captures a wide color distribution (Fig. 4d), which is critical for evaluating color-based anomaly detection approaches. (c) Anomalous Region Proportion: Compared to Real-IAD [39], Real-IAD Variety exhibits a broader and more uniform distribution of anomaly region proportions (Fig. 4a), encompassing defects of varying sizes and thereby increasing detection difficulty. (d) Defect Aspect Ratio: The defect aspect ratio distribution in Real-IAD Variety is comparably diverse to Real-IAD (Fig. 4b), presenting substantial challenges for models to discriminate defects of various geometric configurations. (e) Distribution of Normal and Anomalous Images: Real-IAD Variety contains a substantially more anomalous images than normal images, addressing the “data scarcity” bottleneck in real-world industrial environments (first row of Fig. 5). (f) Defects Spatial and Scale Distribution: The spatial distribution of defect centers (second row in Fig. 5) reveals that anomalies occur across almost all possible image coordinates, demonstrating high spatial diversity. The defect bounding boxes relative to the original resolution shows a high density near the origin (third row in Fig. 5), indicating a predominance of challenging small-scale defects. (g) Anomalous Area Ratio Distribution: The distribution of anomalous regions (fourth row in Fig. 5) approximately follows a normal distribution on a logarithmic scale.
Representative Defect Visualization. Figs. 1 and 3d presents qualitative visualizations of 27 distinct defect types across various object categories. Notably, identical defect types manifest in diverse visual forms across different materials, reflecting the complexity of real-world industrial scenarios. These defect types, derived from authentic production line observations, enhance the practical applicability of Real-IAD Variety.
@@ 4.1 Dataset Protocol and Evaluation Metrics
Training and Testing Protocol. The Real-IAD Variety dataset is partitioned into training and testing subsets, comprising 3,991 normal samples (19,955 images) for training and 35,799 samples (3,990 normal and 31,809 anomalous, totaling 178,995 images) for testing. Notably, the testing subset exhibits a nearly balanced distribution between normal and anomalous images. To systematically evaluate the impact of category count on model performance, we partition Real-IAD Variety into three subsets: S1, S2, and S3, containing 30, 60, and 100 categories, respectively. The category selection employs randomization to mitigate potential biases in color and material representation.
Table 3: Quantitative MUAD performance comparisons on Real-IAD Variety dataset. Dis., Emb., and Rec. represent discrimination-based, embedding-based, and reconstruction-based methods, respectively. The best results are indicated in bold.
Paradigms Methods Backbone I-ROC I-PR I-F1max P-ROC P-PR P-F1max P-PRO
Dis. DRAEM [42] scratch 54.9 89.3 93.1 53.7 9.4 8.8 21.6
SimpleNet [29] WRes-50 53.7 88.5 93.0 62.6 4.8 8.8 29.5
Emb. CFA [25] WRes-50 52.9 88.5 93.0 52.9 2.7 5.6 14.1
CFLOW [14] WRes-50 63.8 91.7 93.1 85.2 14.8 20.0 59.5
Rec. UniAD [41] Eff-B4 67.5 92.4 93.3 87.1 18.0 24.0 62.8
OneNIP [11] Eff-B4 73.3 94.2 93.5 88.3 24.0 29.4 64.5
RD [5] WRes-50 72.8 93.9 93.6 88.9 22.3 28.1 68.9
DeSTSeg [45] Res-18 75.0 95.0 93.2 65.6 37.2 31.2 37.7
MambaAD [17] Res-34 81.7 96.2 93.9 91.2 33.1 37.7 73.5
LGC [9] WRes-50 81.0 96.0 94.0 91.7 31.9 37.0 74.6
Dinomaly [16] ViT-B-14 85.4 97.2 94.5 91.5 42.8 45.8 75.6
Dinomaly+ (Ours) ViT-B-14 87.1 97.3 94.0 91.9 49.7 49.2 76.4
Evaluation Metrics. Following established anomaly detection protocols, we employ seven standard metrics to quantify performance across both image-level anomaly classification and pixel-level anomaly localization tasks. These seven metrics include: Image-level Area Under the Receiver Operating Characteristic curve (I-ROC), Area Under the Precision-Recall curve (I-PR), and Maximum F1-score (I-F1max); alongside their pixel-level counterparts (P-ROC, P-PR, P-F1max) and Pixel-level Per-Region Overlap (P-PRO) for fine-grained defect localization assessment.
@@ 4.2 Benchmark on Multi-Class Unsupervised AD
Unless otherwise specified, MUAD models are trained at a resolution of $\times256$ for 100 epochs. For Dinomaly [16], we follow the original configuration, utilizing a $\times392$ crop from $\times448$ input resolution for 50,000 steps (e.g., approximately 160 epochs for Real-IAD Variety). Additionally, we introduce Dinomaly+, an enhanced version that extends the baseline Dinomaly [16] by integrating a refined segmentation head [11]. This model is fine-tuned for 4 epochs on pseudo-anomalous data to facilitate high-precision, coarse-to-fine anomaly localization.
@@
Competitive Methods. We evaluate state-of-the-art unsupervised anomaly detection methods to comprehensively assess their scalability to large-scale datasets. Specifically, we benchmark discrimination-based methods (DRAEM [42], SimpleNet [29]), embedding-based methods (CFA [25], CFLOW-AD [14]), and reconstruction-based methods (UniAD [41], OneNIP [11], RD [5], DesTSeg [45], MambaAD [17], LGC [9], Dinomaly [16], and Dinomaly+) on Real-IAD Variety. Considering computational resource constraints, we focus on representative and strong baselines, SimpleNet [29], CFLOW-AD [14], UniAD [41], OneNIP [11], MambaAD [17], LGC [9], Dinomaly [16], and Dinomaly+ for comprehensive MUAD evaluation. Notably, PatchCore [33] is excluded from this comparison due to its memory bank mechanism, which incurs prohibitive memory consumption as category count increases, leading to potential out-of-memory issues on large-scale datasets like Real-IAD Variety.
Table 4: Comprehensive comparison of MUAD methods on Real-IAD and Real-IAD Variety datasets. $\Delta$ represents the absolute difference between the maximum and minimum values across Real-IAD Variety scales (S1, S2, S3, and Full) for each method, highlighting performance sensitivity to category expansion, and the same below.
Methods Datasets Classes I-ROC I-PR I-F1max P-ROC P-PR P-F1max P-PRO
SimpleNet [29] Real-IAD 30 54.9 50.6 61.5 76.1 1.9 4.9 42.4
Real-IAD Variety S1 30 75.4 95.2 93.7 83.4 22.1 26.6 52.7
Real-IAD Variety S2 60 67.9 93.2 93.2 80.4 14.4 19.2 46.0
Real-IAD Variety S3 100 61.0 91.0 93.1 78.0 9.2 13.8 41.8
Real-IAD Variety 160 53.7 88.5 93.0 62.6 4.8 8.8 29.5
$\Delta\downarrow$ 21.7 6.7 0.7 20.8 17.3 17.8 23.2
CFLOW [14] Real-IAD 30 77.0 75.8 69.9 94.8 17.6 21.7 80.4
Real-IAD Variety S1 30 77.9 95.4 93.8 87.5 27.6 26.6 63.2
Real-IAD Variety S2 60 72.5 94.1 93.3 86.7 23.6 24.5 61.3
Real-IAD Variety S3 100 72.5 94.1 93.3 86.7 23.6 24.5 61.3
Real-IAD Variety 160 63.8 91.7 93.1 85.2 14.8 20.0 59.5
$\Delta\downarrow$ 14.1 3.7 0.7 2.3 12.8 6.6 3.7
UniAD [41] Real-IAD 30 83.1 81.2 74.5 97.4 23.3 30.9 87.1
Real-IAD Variety S1 30 78.5 95.4 94.2 88.9 29.5 34.1 67.1
Real-IAD Variety S2 60 71.6 93.5 93.7 87.3 21.7 27.6 64.8
Real-IAD Variety S3 100 70.2 93.1 93.5 87.6 19.0 25.0 64.3
Real-IAD Variety 160 67.5 92.4 93.3 87.1 18.0 24.0 62.8
$\Delta\downarrow$ 11.0 3.0 0.9 1.8 11.5 10.1 4.3
OneNIP [11] Real-IAD 30 86.3 84.9 77.2 98.2 37.0 41.5 88.6
Real-IAD Variety S1 30 82.8 96.7 94.4 90.1 36.7 39.1 69.0
Real-IAD Variety S2 60 78.7 95.6 94.0 89.2 30.7 34.9 67.2
Real-IAD Variety S3 100 76.5 94.9 93.7 89.2 27.0 32.2 66.5
Real-IAD Variety 160 73.3 94.2 93.5 88.3 24.0 29.4 64.5
$\Delta\downarrow$ 9.5 2.5 0.9 1.8 12.7 9.7 4.5
MambaAD [17] Real-IAD 30 87.0 85.3 77.6 98.6 32.4 38.1 91.2
Real-IAD Variety S1 30 87.4 97.5 95.0 91.5 40.3 43.2 75.0
Real-IAD Variety S2 60 76.2 94.9 93.8 89.0 25.9 31.5 68.1
Real-IAD Variety S3 100 72.5 93.7 93.6 88.5 21.5 27.5 66.5
Real-IAD Variety 160 81.7 96.2 93.9 91.2 33.1 37.7 73.5
$\Delta\downarrow$ 14.9 3.8 1.4 3.0 18.8 15.7 8.5
LGC [9] Real-IAD 30 85.9 83.0 77.1 98.1 25.3 33.3 90.7
Real-IAD Variety S1 30 84.1 96.7 94.8 91.0 36.1 39.9 73.7
Real-IAD Variety S2 60 82.7 96.4 94.1 91.9 34.5 39.9 76.6
Real-IAD Variety S3 100 82.4 96.2 94.3 91.8 32.8 37.9 75.0
Real-IAD Variety 160 81.0 96.0 94.0 91.7 31.9 37.0 74.6
$\Delta\downarrow$ 3.1 0.7 0.8 0.9 4.2 2.9 2.9
Dinomaly [16] Real-IAD 30 89.3 86.5 80.1 98.9 42.8 47.1 93.9
Real-IAD Variety S1 30 91.4 98.4 95.8 92.7 56.9 56.1 79.8
Real-IAD Variety S2 60 89.7 98.1 95.2 92.6 53.1 53.7 79.7
Real-IAD Variety S3 100 87.9 97.7 94.9 92.3 48.3 50.1 78.2
Real-IAD Variety 160 85.4 97.2 94.5 91.5 42.8 45.8 75.6
$\Delta\downarrow$ 6.0 1.2 1.4 1.2 14.1 10.3 4.2
Dinomaly+ Real-IAD 30 89.5 86.3 80.7 98.8 45.6 48.6 93.1
Real-IAD Variety S1 30 92.4 98.6 96.0 92.7 59.2 56.1 79.3
Real-IAD Variety S2 60 90.8 98.3 95.3 92.9 56.3 54.4 79.6
Real-IAD Variety S3 100 89.4 97.9 94.7 92.4 53.6 52.5 78.4
Real-IAD Variety 160 87.1 97.3 94.0 91.9 49.7 49.2 76.4
$\Delta\downarrow$ 5.3 1.3 2.0 6.9 9.5 6.9 3.2
Results and Analysis under MUAD. Table 3 presents quantitative results of different methods on Real-IAD Variety. Several critical observations emerge that differ from findings on small-scale datasets. (1) All evaluated methods experience substantial and non-saturating performance degradation on Real-IAD Variety, indicating the dataset’s significant challenge for future research and confirming successful avoidance of the metric saturation phenomenon frequently observed in smaller-scale IAD benchmarks; (2) Discrimination-based approaches (DRAEM [42], SimpleNet [29]) and the embedding-based method CFA [25] exhibit near-failure performance, with I-ROC scores approaching the 50% random guessing threshold. This demonstrates their severe lack of generalization capability when confronted with large-scale, high-variance normal data distributions; (3) Reconstruction-based methods maintain relatively satisfactory performance despite the challenging conditions. Notably, Dinomaly+ (building upon Dinomaly [16] and OneNIP [11]) achieves the best overall results across most metrics (e.g., 87.1% I-ROC and 49.7% P-PR), underscoring its superior generalization and robustness across large-scale multi-class scenarios. In contrast, UniAD [41], an early MUAD method, exhibits one of the most severe performance drops among reconstruction-based approaches.
@@
Performance Trends with Increasing Categories. To evaluate performance trends as category count increases, we selected eight representative methods across different paradigms. Table 4 presents results on Real-IAD and Real-IAD Variety with incrementally increasing categories: S1 (30 classes), S2 (60 classes), S3 (100 classes), and the full 160 classes (visualized in Fig. 6).
@@
Several consistent conclusions emerge under the MUAD setting. (1) Real-IAD Variety poses a more significant challenge than the original Real-IAD. Even at the same category scale (30 classes), the increased intra-class variance and higher resolution complexity in Real-IAD Variety lead to consistent performance drops across most baselines. (2) A fundamental scalability challenge of MUAD is revealed: consistent performance degradation occurs across all evaluated methods as category count increases. Crucially, the magnitude of degradation is method-dependent, highlighting that different architectures possess varying levels of category generalization capability and robustness to scale-up. (3) Methods leveraging powerful backbones (e.g., DINOv2-R [4]) and advanced architectures (e.g., Transformer decoders), such as Dinomaly and Dinomaly+, consistently achieve substantial performance superiority across all datasets. This advantage is especially prominent on high-scale benchmarks, underscoring their effectiveness over normalizing flow-based (e.g., CFLOW-AD) or simpler backbone-based models (e.g., SimpleNet, UniAD, MambaAD).
The performance degradation of MUAD methods when scaling to a large number of categories primarily stems from two factors. First, as the number of categories increases, models often suffer from a capacity-diversity conflict, converging toward a “generic” feature space that loses category-specific nuances essential for spotting subtle anomalies. To address this, recent methods such as OneNIP [11] and LGC [9] introduce category-level information via explicit visual prompts or class-aware contrastive learning to resolve feature reconstruction and preserve class-specific details. Second, reconstruction-based multi-class models are susceptible to identity mapping. In these cases, the model’s excessive generalization capability allows it to reconstruct the anomalies themselves rather than restoring them to a normal state, thereby failing to produce a significant residual for detection. The state-of-the-art Dinomaly [16] partially mitigates this by leveraging a more powerful vision encoder (e.g., DINOv2-R) and simple dropout to achieve more robust performance.
Table 5: Quantitative comparison of ZSAD and FSAD methods on Real-IAD and Real-IAD Variety datasets. Results for FSAD methods represent the average and standard deviation over three runs with random seeds.
Methods Datasets Classes shot I-ROC I-PR I-F1max P-ROC P-PR P-F1max P-PRO
AnomalyCLIP [49] Real-IAD 30 0 69.1 63.5 65.4 95.2 26.8 34.7 84.9
Real-IAD Variety S1 30 0 68.6 88.9 88.6 88.7 36.5 40.3 71.7
Real-IAD Variety S2 60 0 68.8 88.6 88.3 88.5 36.1 40.0 72.4
Real-IAD Variety S3 100 0 69.5 88.9 88.4 89.5 36.2 40.2 74.2
Real-IAD Variety 160 0 69.3 88.9 88.3 89.3 35.3 39.6 74.6
$\Delta\downarrow$ 0.9 0.3 0.3 1.0 1.2 0.7 2.9
AdaCLIP [2] Real-IAD 30 0 71.1 67.2 67.0 95.6 33.6 38.1 90.2
Real-IAD Variety S1 30 0 69.7 89.3 88.7 87.9 30.8 35.4 70.0
Real-IAD Variety S2 60 0 70.1 88.9 88.7 88.0 32.0 36.6 71.3
Real-IAD Variety S3 100 0 71.3 89.4 88.6 88.9 32.8 37.0 72.7
Real-IAD Variety 160 0 71.1 89.3 88.6 89.1 32.6 36.8 73.1
$\Delta\downarrow$ 1.6 0.5 0.1 1.2 2.0 1.6 3.1
VCPCLIP [31] Real-IAD 30 0 72.5 70.9 68.2 95.9 28.8 34.5 83.7
Real-IAD Variety S1 30 0 71.4 90.0 88.7 87.3 34.9 38.1 67.2
Real-IAD Variety S2 60 0 73.0 90.4 88.7 87.9 37.5 40.3 69.5
Real-IAD Variety S3 100 0 72.9 90.5 88.6 89.1 37.6 40.4 71.1
Real-IAD Variety 160 0 72.6 90.2 88.6 89.3 36.9 39.9 70.9
$\Delta\downarrow$ 1.6 0.5 0.1 2.0 2.7 2.3 3.9
AdaptCLIP [10] Real-IAD 30 0 74.2 70.8 68.9 94.9 27.9 35.4 83.1
Real-IAD Variety S1 30 0 73.2 94.7 93.7 88.0 36.5 40.0 68.3
Real-IAD Variety S2 60 0 72.4 94.2 93.5 88.1 36.3 39.9 69.2
Real-IAD Variety S3 100 0 72.6 94.2 93.4 89.0 36.6 40.1 71.1
Real-IAD Variety 160 0 72.9 94.3 93.4 89.1 36.1 39.9 70.8
$\Delta\downarrow$ 0.4 0.5 0.3 1.1 0.5 0.2 2.8
[…]
$\Delta\downarrow$ 1.9 0.3 0.2 1.5 2.3 1.2 3.2
[…]
$\Delta\downarrow$ 1.4 0.3 0.2 1.4 1.2 1.0 3.6
[…]
$\Delta\downarrow$ 0.8 0.2 0.2 1.3 0.7 0.7 3.0